OpenStudy (anonymous):

weighing the baby at the clinic was a problem. the baby would not keep still and caused the scales to wo wobble. so i held the baby and stepped on the scales while the nurse read off 78kg. then the nurse held the baby while i read off 69 kg. finally, i held the nurse while the baby read off 137kg. what is the combined of all three? weighing the baby at the clinic was a problem. the baby would not keep still and caused the scales to wo wobble. so i held the baby and stepped on the scales while the nurse read off 78kg. then the nurse held the baby while i read off 69 kg. finally, i held the nurse while the baby read off 137kg. what is the combined of all three? @Mathematics

OpenStudy (stormfire1):

You only need to find the baby's weight since you know yours + the nurse's is 137kg. Set up some simultaneous equations: let x = you let y = nurse let z = baby x + z = 78 y + z = 69 x + y = 137 Rewrite x = 78 - z y = 137 - x z = 69 - y Substitution the equations for z and y into the first equation so you can solve for x: x = 78 - (69 - (137 - x)) x = 78 - (69 - 137 + x) x = 78 - 69 + 137 - x 2x = 146 x = 73 Find the baby's weight based on the earlier equation: 78 - 73 = 5 Since you know you and the nurse weigh 137kg, just add the baby's weight to get a total of 142 kg.